The ratio of the measures of a pair of sides of a rectangle is 7:12. If the perimeter of the rectangle is 95, what is the length of each side? Answer by mananth(16949) 
 The ratio of the measures of a pair of sides of a rectangle is 7:12. If the perimeter of the rectangle is 95, what is the length of each side?
7/12= ratio of adjacent sides
Perimeter = 95
Sum of adjacent sides = 95/2 = 47.5
7/12= ratio of adjacent sides
7+12=19
19 parts = 47.5
1part = 2.5
7 parts = 17.5
12 parts = 30
The length & breadth are 17.5 & 30